OpenStudy (anonymous):

the unit digit in the product (122)^173 is

OpenStudy (anonymous):

\[ 2^1=2\\ 2^2=4\\ 2^3=8\\ 2^4=16 (6)\\ 2^5=32 (2)\\ 2^6=64 (4)\\ 2^7=128 (8)\\ 2^8=256 (6)\\ 2^9=512 (2)\\ 2^{10}=1024 (4)\\ \]The pattern here is \(2,4,8, 6,2,4,8,6,\ldots\)

OpenStudy (anonymous):

Figure out the remainder of \(173\).

OpenStudy (anonymous):

The remainder when dividing by \(4\).

OpenStudy (luigi0210):

\[\Huge \mathbb {\color{red}{WELCOME}~~\color{blue}{TO}~~\color{green}{OPENSTUDY} \color{purple}{\ddot\smile}}\]

OpenStudy (anonymous):

8

OpenStudy (anonymous):

8 is the remainder ah

OpenStudy (anonymous):

No, the remainder after dividing by 4. It should be between 0, 1, 2, and 3.

OpenStudy (anonymous):

why divied by 4

OpenStudy (anonymous):

It's the remainder.

OpenStudy (anonymous):

Starting at 173 We can remove 100 since it's divisible by 100 73 We can remove 40 33 We can remove 20 13 We can remove 12 1 So the remainder is 1

OpenStudy (anonymous):

\(2^1=2\) so the unit digit will be 2.